OpenStudy (anonymous):

okay I know you said you leave the square root alone but I don't think so on this one (see pic) @johnweldon1993

OpenStudy (johnweldon1993):

lol okay show me :)

OpenStudy (anonymous):

\[\sqrt{8}+\sqrt{50}\]

OpenStudy (johnweldon1993):

Ahh yes...well this is another case :) Well lets see....what are all the numbers that multiply to make 8? 1, 2, 4, and 8 right? Lets ask ourselves...which of these is a perfect square? hmm well 4 is 2² so lets use that \[\sqrt{8} = \sqrt{2}\sqrt{4}\] See what I did there? And now we can simplify that because we know √4 is 2 \[\sqrt{8} = 2\sqrt{2}\] We can do this to the other one as well...50...what 2 numbers *making 1 a perfect square* multiply to make 50?

OpenStudy (johnweldon1993):

I hope that made sense?? Here just let me continue and you can ask if theres a part thats confusing \[\sqrt{8} + \sqrt{50}\] can become \[\sqrt{2}\sqrt{4} + \sqrt{2}\sqrt{25}\] This is because we know that √2 times √4 = √8.....and √2 times √25 = √50 Now this is helpful for us...because we know that 4 is a perfect square of 2.....and we know that 25 is a perfect square of 5 *2² = 4 and 5² = 25* so now when we simplify this...we have \[2\sqrt{2} + 5\sqrt{2}\] Now we have add them like the questions before...

OpenStudy (anonymous):

25 is one and.........(thinking)......uhmmm.......i cant think of another one:P sorry

OpenStudy (johnweldon1993):

Ahh yeah see that is what I was going for! lol....25 times 2 = 50....and we only need 1 number to be a perfect square...which is the one that we needed here 25 = 5².....good job...now refer to the above post :)

OpenStudy (anonymous):

\[7\sqrt{2}\] would this be my total answer?? @johnweldon1993

OpenStudy (johnweldon1993):

That would be correct! :)